OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4.

OpenStudy (anonymous):

Here we have a point, and a hint at slope. The point is (2,6) and the slope parallel to the given line is the same slope of the line The slope PERPENDICULAR to the line is the OPPOSITE RECIPROCAL ( flip the fraction upside down, if it is a whole number put 1 over the whole number) Since your given equation is y = -3x - 4, the slope of this line is -3. Since we are using the PARALLEL slope to this line, we use the EXACT slope -3. (Note if we have perpendicular, we would use slope 1/3) Now that you have slope m = -3 and point (2,6) Plug this into the POINT SLOPE form y-y1 = m(x-x1)